Operation Summary for January 25, 2010 - January 27, 2010 |
Monday Jan. 25 The day shift began with Operations monitoring the Tevatron (TeV) store 7534 (59.01E30) and the stash (260.6E10), with the Antiproton Source stacking (15.0E10), with Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Linac personnel tuned the H- Source to allow it to run for another day. Operations terminated store 7534 at 8:30 AM and then put the TeV into HEP shot setup. Operations established store 7535 at 10:15 AM with an average initial luminosity (AIL) of 206.9E30. During the evening shift, at 7:05 PM the Pelletron tripped off; a Recycler person came in and fixed it. Operations disabled the MTest beam at 10 PM. Tuesday Jan. 26 The midnight shift began with Operations monitoring store 7535 (57.93E30) and the stash (331.5E10), with Pbar stacking (15.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7535 at 2 AM, put the TeV into shot setup, and also turned off the beam to Pbar. Operations established store 7536 at 4 AM with an AIL of 244.89E30. Operations turned off the beam to NuMI and MiniBooNE at 5:30 AM in preparation for the shutdown to repair the H- Source. During the day shift, Operations maintained a small Recycler stash throughout the day for studies. The rest of the complex was in shutdown for various safety system tests and maintenance work while the Linac personnel removed the H- Source for cleaning and replaced a failing ion pump. During the evening shift, Interlock technicians completed the Booster safety system tests. TeV personnel conducted crystal collimator studies. Wednesday Jan. 27 The midnight shift began with the H- Source down for repairs, with Operations monitoring store 7536 (35.77E30), with Recycler personnel conducting studies, and the rest of the complex off. During the midnight shift, the Recycler studies finished at 4 AM and safety system tests began. Operations had the MI and Recycler ready for access by 5 AM. Operations terminated store 7536 at 5:30 AM and then made TeV, CDF, and D0 ready for access. The Plans for Wednesday The plans for today are to finish up maintenance work. The Preaccelerator might be ready to resume delivering beam on Thursday (1/28/10). |
|
More Information Comments and Suggestions |
| last modified 02/03/2010 email Fermilab |
| Security, Privacy, Legal | |